FOURIER TRANSFORM SPECTRA AND INVERTED TORSIONAL STRUCTURE FOR A CH_{3}-BENDING FUNDAMENTAL OF CH_{3}OH

\underline{R.M. LEES}, \underline{LI-HONG XU} and ANNA K. KRISTOFFERSON, Department of Physical Sciences, University of New Brunswick, Saint John, NB, Canada E2L 4L5; MICHAEL LOCK and B.P. WIN-NEWISSER, Physikalisch-Chemisches Institut, Justus Liebig University, D-35392 Giessen, Germany; J.W.C. JOHNS, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, ON, Canada K1A 0R6.

The high-resolution Fourier transform spectrum of CH_{3}OH has been investigated in the 1400-1650 cm^{−1} region, containing the CH_{3}-bending fundamental bands. Twenty-two perpendicular ΔK = +1 subbands have been identified so far, with origins ranging from 1490 to 1570 cm^{−1} for transitions K = 2 ← 1 up to 11 ← 10 for various torsional symmetries. Assignment of the subbands to the ν_{4} in-plane (A′) or ν_{10} out-of-plane (A′′) asymmetric methyl-bending modes is not yet clear, but the one subband so far observed with resolved K-doublet structure suggests c-type selection rules consistent with a ν_{10} vibrational assignment. The pattern of the K-reduced torsion-vibration energy τ-curves is inverted compared to the normal 1-dimensional picture for n = 0 torsional levels, in agreement with prediction based on fitting torsional variation of ab initio CH_{3}-bending frequencies to a local mode model. However, the periodicity of the curves is unusual and significantly different from the ground state. The vibrational energy for the bending mode is 1481 cm^{−1}, and the mean B-value is 0.008 cm^{−1} higher than that of the vibrational ground state.